OpenStudy (wesdg1978):

Use definition: f(x) = lim h→0 f(x + h) − f(x)/h to find f ' (x) and then find the tangent line to the graph of y=f(x) at x=a f(x)=sqrt(x+1); a=8 I'm not sure what to do here.

OpenStudy (anonymous):

plug ins sqrt(x+1) to your f(x) definition f(x) = lim h→0 f(x + h) − f(x)/h --- or it could be easier to plug in a from the start plug ins sqrt(8+1) to your f(x) definition f(sqrt(8+1)) = lim h→0 f((sqrt(8+1)) + h) − f(sqrt(8+1))/h

OpenStudy (amistre64):

thanks to the sqrt, conjugates are involved

OpenStudy (wesdg1978):

OK, the answer is supposed to be 1/(2sqrt(x+1)) and y=1/6x+5/3 I'm not seeing how to get there.

OpenStudy (amistre64):

define the starting setup apply conjugates simplify out the h dividor

OpenStudy (amistre64):

cancel out the h dividor? .... simplify, cancel .... either way it gets taken care of

OpenStudy (anonymous):

I think i did a mistake in mines.

OpenStudy (amistre64):

\[\lim~\frac{f(x+h)-f(x)}{h}\] \[\lim~\frac{f(x+h)-f(x)}{h}\frac{f(x+h)+f(x)}{f(x+h)+f(x)}\] \[\lim~\frac{f^2(x+h)-f^2(x)}{h(f(x+h)+f(x))}\] with the given f(x) function, that undoes the sqrts and you can factor out an h above

OpenStudy (amistre64):

as such, as h goes to 0\[\lim~\frac{1}{f(x+0)+f(x)}=\frac{1}{2f(x)}\]

OpenStudy (wesdg1978):

OK, that makes sense, now I need to get the tangent line to the graph at x=a and a=8. How do I do that.

OpenStudy (amistre64):

a line equation is from algebra: the tangent line form will with be point slope, or slope intercept ... not that it really matter; the results are the same regardless y = f'(a)(x-a) + f(a)

OpenStudy (amistre64):

once you know f', the defines the slope for any given a

OpenStudy (wesdg1978):

So, y=1/(2sqrt(x+1)) (x-8)+f(sqrt8+1) ?

OpenStudy (amistre64):

close, define f'(a) not f'(x), and dont use f(sqrt(x+1)) thats nowhere defined .. other than being a typo f(x) = sqrt(x+1) y= 1/(2 sqrt(8+1)) (x-8) + sqrt(8+1) ?

OpenStudy (wesdg1978):

So this is what I got from that: y=1/6x-8+3 or y=1/6x-5 The answer in the back of the book is y=1/6x+5/3

OpenStudy (amistre64):

\[y=\frac{1}{2\sqrt9}(x-8)+\sqrt{9}\] \[y=\frac{1}{2.3}(x-8)+3\] \[y=\frac{1}{6}(x-8)+3\] \[y=\frac{1}{6}x-\frac 86+3\] \[y=\frac{1}{6}x-\frac 43+3\] \[y=\frac{1}{6}x-\frac 43+\frac 93\] \[y=\frac{1}{6}x+\frac 53\]

OpenStudy (wesdg1978):

I just wasn't paying close enough attention, thanks for your help!

OpenStudy (amistre64):

it happens ;) good luck